this is my question
how do i write an equation of the line through the given point with the given slope? (0,2);m=4/5 then write in slope intercept form?

Answer by prabhjyot(165)

The line through (0,2) with slope 4/5 has equation
y = y1 + m(x - x1)
= 2 + 4/5(x - 0)
= 2 + (4/5)x
= 4/5x + 2